OpenStudy (anonymous):

find dy/du, du/dx and dy/dx y= u^2 u=4x+7

myininaya (myininaya):

dy/du=2u du/dx=4 dy/dx=2*u*4=8u=8(4x+7) dy/dx=dy/du*du/dx (chain rule)